# Concrete examples of the rate of convergence of Chernoff approximations: numerical results for the heat semigroup and open questions on them (with appendix: full list of pictures and Python code)

### v4

@@ 1 Introduction
$\left\{\begin{array}[]{ll}u^{\prime}_{t}(t,x)=u_{xx}^{\prime\prime}(t,x)\textrm{ for }t>0,x\in\mathbb{R}^{1}\\ u(0,x)=u_{0}(x)\textrm{ for }x\in\mathbb{R}^{1}\end{array}\right$. (1)
@@ 3.4.3 Infinitely smooth initial conditions
3.5 Unexplained behavoir of error was found in simulation
We use the context of Remark 6 on p. 6 and the notation introduced in Remark 10 on p. 10. There are no known theoretical bounds on the rate of convergence for initial conditions that do not belong to the domain of the generator of the semigroup. For the heat semigroup, this means that initial condition do not have two uniformly continuous bounded derivatives. We studied this numerically, so the results we found are the only evidence known at the moment. We have collected the results of our computer simulation in the tables above.
@@ 3.5 Unexplained behavoir of error was found in simulation
Generally we see that $\alpha_{S}<\alpha_{G}$ in all cases (all initial conditions that we have tested), which means that a higher order of the Chernoff tangency of the Chernoff function to the generator of the semigroup provides a better order of the Chernoff approximation error decrease. Also, we see that a higher smoothness of the initial condition provides a better order of decreasing of the Chernoff approximation error for both Chernoff functions $S$ and $G$ and for smoothness $\xi\in(0,1]$. These observations are not surprising in the context of theorem 4.
Let us now list the observations that were unexpected and are not predicted by theorem 4. These are effects that we see numerically but do not know how to explain theoretically.
@@
  2. 2.
    Linear approximation of the error in logarithm scale is not equally perfect. For smooth $u_{0}(x)=\sin(x)$ we see that points fit the line perfectly (fig. 1.2), as well as for non-smooth $u_{0}(x)=|\sin(x)|^{1/4}$, see fig. 3.2. But for smooth $u_{0}(x)=e^{-x^{4}}$ we see that points do not fit the line well (fig. 6.2). It seems that a higher smoothness provides a worse line fit, but for $u_{0}(x)=\sin(x)$ the fit is very good.
@@
    The order of error behaves in the opposite to the free term of the error. In other words: $\alpha_{S}<\alpha_{G}$ for all initial conditions, but $\beta_{S}>\beta_{G}$ for all initial conditions except $u_{0}(x)=\sin(x)$ (large difference in $\beta$) and $u_{0}(x)=e^{-|x|^{5/2}}$ (small difference in $\beta$).
    Smoothness matters, but what else is important? Two examples are provided below.
@@
    The initial conditions $\sin(x)$, $e^{-x^{2}}$, $e^{-x^{4}}$ are infinitely smooth, but provide a very different behavior of the error.
    For $e^{-x^{2}}$ the regression line parameters are not included in the tables to avoid making two equally colored lines corresponding to infinite smoothness in figures 11 and 12. For $e^{-x^{2}}$ we have $\alpha_{S}=-2.1028$, $\beta_{S}=-3.2967$, $\alpha_{G}=-1.0177$, $\beta_{G}=-3.3874$. Graphs for $e^{-x^{2}}$ are provided below on figure 14; why they are so different from graphs for $e^{-x^{2}}$ (fig. 5.2) and for $\sin(x)$ (fig. 1.2)?
    For $\sin(x)$ the total error for the function $S$ is smaller than for $G$, starting from $n=1$, but for $e^{-x^{2}}$ and $e^{-x^{4}}$ this holds only starting form larger values of $n$. For $\sin(x)$ and $e^{-x^{2}}$ $\alpha_{S}\approx-2$, it is in line with the fact that $S$ has the second order of Chernoff tangency to the generator of the semigroup. But for $e^{-x^{4}}$ unexpectedly we have the order $\alpha_{S}\approx-3$, and we do not know why.
@@
    The initial conditions $\sin(x)\cdot|\sin(x)|$ and $\sin(x)\cdot|\sin(x)|+e^{-x^{2}}$ have the same smoothness and provide a similar behavior of the error. Graphs and values are provided below.
    fig. 19, $1\leq n\leq 11$, $u_{0}(x)=\sin(x)\cdot|\sin(x)|$, $t=\frac{1}{2}$
    For $u_{0}(x)=\sin(x)\cdot|\sin(x)|$, $t=\frac{1}{2}$ we have $\alpha_{S}=-1.2573$, $\beta_{S}=-2.6904$, $\alpha_{G}=-1.1082$, $\beta_{G}=-3.1804$.
    fig. 20, $1\leq n\leq 11$, $u_{0}(x)=\sin(x)\cdot|\sin(x)|+e^{-x^{2}}$, $t=\frac{1}{2}$
    For $u_{0}(x)=\sin(x)\cdot|\sin(x)|+e^{-x^{2}}$ we have $\alpha_{S}=-1.3242$, $\beta_{S}=-2.4434$, $\alpha_{G}=-0.9844$, $\beta_{G}=-2.8821$.
@@
    The initial conditions $u_{0}=\sin(x)$ and $u_{0}=\cos(x)$ provide the same values of $\alpha$ and $\beta$ for both Chernoff functions $S$ and $G$.
4 Conclusion
We studied a nontrivial concrete case of a $C_{0}$-semigroup – the heat semigroup, and numerically tested two known Chernoff functions for it. Our aim was to collect numerical observations on the speed of vanishing of the error of Chernoff approximations. We found some interesting effects with the use of numerical simulation; we described these effects but have no theoretical explanation for them. So, research in this direction is far from ending and opens up new challenges.

### v5

@@ 1 Introduction
$\left\{\begin{array}[]{ll}u_{t}(t,x)=u_{xx}(t,x)\textrm{ for }t>0,x\in\mathbb{R}^{1}\\ u(0,x)=u_{0}(x)\textrm{ for }x\in\mathbb{R}^{1}\end{array}\right$. (1)
@@ 3.4.3 Infinitely smooth initial conditions
4 Conclusion
Let us recall the context of what we did and point out the key findings. $C_{0}$-semigroups have various applications in mathematics and mathematical physics [6, 7]. In the case where the infinitesimal generator $L$ of the semigroup $(e^{tL})_{t\geq 0}$ is a differential operator with variable coefficients, often Chernoff’s theorem is the most clear way to express $e^{tL}$ explicitly in terms of $L$, e.g., in terms of $a,b,c$ if $(Lf)(x)=a(x)f^{\prime\prime}(x)+b(x)f^{\prime}(x)+c(x)f(x)$, $x\in\mathbb{R}$. In this approach we have $e^{tL}=\lim_{n\to\infty}C(t/n)^{n}$ where $C$ is the so-called Chernoff function for the operator $L$, and $C(t/n)^{n}$ is the composition of $n$ copies of linear bounded operator $C(t/n)$. Expressions $C(t/n)^{n}$ are called the Chernoff approximations to $e^{tL}$ provided by Chernoff function $C$, and it is possible [5, 21, 22, 35] to express $C$ as a short formula containing $a,b,c$ as parameters. One should keep in mind that for a given operator $L$ there exists only one semigroup $e^{tL}$ but many Chernoff functions $C$, and they provide different rate of convergence $C(t/n)^{n}\to e^{tL}$ as $n\to\infty$.
The knowledge of $e^{tL}=\lim_{n\to\infty}C(t/n)^{n}$ allows to represent the bounded solution of the Cauchy problem $u_{t}(t,x)=a(x)u_{xx}+b(x)u_{x}+c(x)u$, $u(0,x)=u_{0}(x)$ as a formula in terms of $a,b,c,u_{0}$, because due to the general theory of $C_{0}$-semigroups [6, 7] we have $u(t,x)=(e^{tL}u_{0})(x)$. Also, one can represent (see [25]) the bounded solution of the equation $a(x)f^{\prime\prime}(x)+b(x)f^{\prime}(x)+(c(x)-\lambda)f(x)=-g(x)$ as a formula in terms of $a,b,c,g,\lambda$ because (again due to the general theory of $C_{0}$-semigroups [6, 7]) we have $f(x)=\int_{0}^{\infty}e^{-\lambda t}(e^{tL}g)(x)dt$. Chernoff’s theorem has other applications, but these two are the most easy to explain in a few words.
Our goal is to study the rate of convergence in the Chernoff theorem numerically (theoretical knowledge in this area is still very limited), and we selected the heat semigroup on the real line as a simple yet informative model case. We rewrite the initial value problem
$\left\{\begin{array}[]{ll}u_{t}(t,x)=u_{xx}(t,x)\textrm{ for }t>0,x\in\mathbb{R}^{1}\\ u(0,x)=u_{0}(x)\textrm{ for }x\in\mathbb{R}^{1}\end{array}\right$.
as $U^{\prime}(t)=LU(t),\,U(0)=u_{0}]$ with $Lf=f^{\prime\prime}$, $u(t,\cdot)=U(t)=e^{tL}u_{0}$. We know that $(e^{tL})_{t\geq 0}$ is a $C_{0}$-semigroup in the Banach space $\mathcal{F}=UC_{b}(\mathbb{R})$ of all uniformly continuous and bounded functions $f\colon\mathbb{R}\to\mathbb{R}$, $\|f\|=\sup_{x\in\mathbb{R}}|f(x)|$. For each $t\geq 0$ we introduce linear bounded operators $G(t)$ and $S(t)$ in $\mathcal{F}$ as
$(G(t)f)(x)=\frac{1}{2}f(x)+\frac{1}{4}f(x+2\sqrt{t})+\frac{1}{4}f(x-2\sqrt{t}),$
$(S(t)f)(x)=\frac{2}{3}f(x)+\frac{1}{6}f(x+\sqrt{6t})+\frac{1}{6}f(x-\sqrt{6t}).$
It is known that $G$ and $S$ are Chernoff functions for the operator $L=\partial^{2}$. Paul Chernoff’s theorem [4] implies that for each $t>0$, each $u_{0}\in\mathcal{F}$ sequences
$d_{n}^{G}=\|e^{tL}u_{0}-G(t/n)^{n}u_{0}\|,\quad d_{n}^{S}=\|e^{tL}u_{0}-S(t/n)^{n}u_{0}\|$
converge to 0 as $n\to\infty$, but it says nothing about the rate of convergence. We try to say at least something via numerical simulation.
We considered 25 initial conditions of different smoothness: trigonometric ($u_{0}(x)=|\sin x|^{\xi}$), exponential ($u_{0}(x)=\exp(-|x|^{\xi}$), and also several infinitely smooth functions ($u_{0}(x)=e^{-x^{2k}}$, $u_{0}(x)=e^{-x^{2}}\sin(x)$). For all of them, we numerically calculated
$\ln d_{n}^{G}\approx\alpha_{G}\ln n+\beta_{G},\quad\ln d_{n}^{S}\approx\alpha_{S}\ln n+\beta_{S}.$
We call $\alpha$ the order of decreasing error, $\beta$ the free term of decreasing error. Function $G$ has the first order of Chernoff tangency to $\partial^{2}$, and function $S$ has the second order of Chernoff tangency to $\partial^{2}$. So, one can expect $\alpha_{G}\approx-1$, $\alpha_{S}\approx-2$, but the numerical experiment shows that the reality is not that simple.
These are not surprising findings that were predicted by the existing theory:
@@ 4 Conclusion
  • We see that $\alpha_{S}<\alpha_{G}$ for all initial conditions $u_{0}$ that we have tested. This means that a higher order of the Chernoff tangency of the Chernoff function to the generator of the semigroup provides a better order of decreasing error. Indeed, blue points are below green points on ($\xi,\alpha$) plane for trigonometric initial conditions (fig. 9) and for exponential initial conditions (fig. 11).
  • The difference between the errors $d_{n}^{G}$ and $d_{n}^{S}$ is obvious if $\xi\in[2,4.5]$, i.e. when $u_{0}\in D(L)$.
  • If $u_{0}\notin D(L)$, i.e. $\xi\in(0,2)$, then this difference is small.
These are effects that we see numerically but do not know how to explain theoretically:
@@
    The linear approximation of error $\ln d_{n}$ vs. $\ln n$ is not equally perfect, and it is often different for different initial conditions. But for $u_{0}(x)=\sin(x)$, $u_{0}(x)=\cos(x)$ the error is identical. For smooth $u_{0}(x)=\sin(x)$ we see that points fit the line perfectly (fig. 1.2), as well as for non-smooth $u_{0}(x)=|\sin(x)|^{1/4}$, see fig. 3.2. But for smooth $u_{0}(x)=e^{-x^{4}}$ we see that points do not fit the line well (fig. 14.2). It seems that a higher smoothness provides a worse line fit, but for $u_{0}(x)=\sin(x)$ the fit is very good.
  2. 2.
@@
    The order of error behaves in the opposite to the free term of the error. In other words: $\alpha_{S}<\alpha_{G}$ for all initial conditions, but $\beta_{S}>\beta_{G}$ (green points are below blue points) for almost all initial conditions, exeptions are $u_{0}(x)=\sin(x)$ (large difference in $\beta$) and $u_{0}(x)=e^{-|x|^{5/2}}$ (small difference in $\beta$). Cf. fig. 9 vs. fig. 10, fig. 11 vs. fig. 12.
@@
    For $\xi\in(0,1)$ dependence of $\alpha_{S}$ and $\alpha_{G}$ on $\xi$ seems to be linear for both trigonometric and exponential $u_{0}$.
    Smoothness matters, but it is clear that something else is also important. Indeed:
@@
    Infinitely smooth initial conditions can provide a different behavior of the error:
    $u_{0}(x)=\sin(x):\ \alpha_{G}\approx-1.0,\alpha_{S}\approx-2.0$
    $u_{0}(x)=e^{-x^{2}}:\ \alpha_{G}\approx-1.0,\alpha_{S}\approx-2.1$
    $u_{0}(x)=e^{-x^{4}}:\ \alpha_{G}\approx-1.4$, $\alpha_{S}\approx-3.1$
    $u_{0}(x)=e^{-x^{6}}:\ \alpha_{G}\approx-1.8,\alpha_{S}\approx-1.8$
@@
    For trigonometric initial conditions $u_{0}$ with $\xi\in[1,4.5]$ the dependence of $\alpha_{S}$ on $\xi$ seems to be linear, but it seems that there is no dependence of $\alpha_{G}$ on $\xi$. Meanwhile, for exponential $u_{0}$ with $\xi\in[1,4.5]$ it seems that there is no dependence of both $\alpha_{G}$ and $\alpha_{S}$ on $\xi$.